# DSSYK at Infinite Temperature: The Flat-Space Limit and the ‘t Hooft Model

### v1

@@ 6.1 The Phase Boundary and the Stretched Horizon
where $x^{i}$ are coordinates parameterizing the $(D-2)$-dimensional plane of the horizon.
@@
Let us introduce a mathematical $(D-1)$-dimensional time-like surface of fixed $\rho$ at the value of $\rho$ for which
@@
To a high approximation the phase boundary is the end of the world. The outer confined region is almost completely empty; the vast majority of degrees of freedom cannot escape the hot region. Only the light $SU(N)$ singlets can escape. When a quark from the hot region hits the phase boundary it is reflected back with a probability very close to $1$ With a small probability it passes through the boundary dragging an antiquark with it, the two forming a meson. A similar thing can be said for gluons. In other words to quote [12] “almost everything is confined” (to the hot plasma-like stretched horizon) [^11] The word confined has two different meanings in this context. Non-singlet degrees of freedom are confined to the deconfined (green) region. In the confined region quarks and gluons are confined to form singlet hadrons..
@@ 6.2 Escape from the Stretched Horizon
The spectrum of singlets is very sparse and unlike the spectrum of quarks (or quarks and gluons) it does not grow with increasing $N.$ This resolves a certain tension noted earlier in [12]. The total number of degrees of freedom is $N$ as seen from the fact that the entropy of ${\rm DSSYK_{\infty}}$ is exactly
$N\log{2}$
@@ 7 Conclusions
A holographic duality should be more than just a claim of a bulk geometry; it should also determine the matter fields and particles which can propagate in the bulk as well as the interactions. Ideally a holographic description of our own universe would include the standard model, neutrino masses, and other features of particle physics. We have taken up the challenge of determining the bulk theory for a particular example: the holographic duality between double-scaled SYK and JT-de Sitter space. The conjecture that ${\rm DSSYK_{\infty}}$ is dual to JT-de Sitter is not new but the identification of the bulk matter sector is. We’ve given evidence that the bulk theory is the ultra-strongly coupled fixed $\bar{g}$ limit of $(1+1)$ dimensional QCD—a string theory with mesons lying on a single Regge trajectory. Unlike the limit $\bar{g}^{2}N$ held fixed, the fixed $\bar{g}$ limit is a theory of interacting meson-strings with $\bar{g}$ being the open string coupling constant. On the ${\rm DSSYK_{\infty}}$ side the role of $\bar{g}$ is played by $\lambda.$ In the limit $\lambda\to 0$ the standard ‘t Hooft model of non-interacting mesons is retrieved.

### v3

@@ 6.1 The Phase Boundary and the Stretched Horizon
where $x^{i}$ are coordinates parameterizing the (D2)(D-2)-dimensional plane of the horizon.
@@
Let us introduce a mathematical (D1)(D-1)-dimensional time-like surface of fixed $\rho$ at the value of $\rho$ for which
@@
To a high approximation the phase boundary is the end of the world. The outer confined region is almost completely empty; the vast majority of degrees of freedom cannot escape the hot region. Only the light $SU(N)$ singlets can escape. When a quark from the hot region hits the phase boundary it is reflected back with a probability very close to $1.$ With a small probability it passes through the boundary dragging an antiquark with it, the two forming a meson. A similar thing can be said for gluons. In other words to quote [12] “almost everything is confined” (to the hot plasma-like stretched horizon) [^15] The word confined has two different meanings in this context. Non-singlet degrees of freedom are confined to the deconfined (green) region. In the confined region quarks and gluons are confined to form singlet hadrons..
@@ 6.2 Escape from the Stretched Horizon
The spectrum of singlets is very sparse and unlike the spectrum of quarks (or quarks and gluons) it does not grow with increasing $N.$ This resolves a certain tension noted earlier in [12]. In the present $D=2$ case the total number of degrees of freedom is $N$ as seen from the fact that the entropy of $\rm DSSYK_{\infty}$} is exactly
$N\log{2},$
@@ 7 Conclusions
A holographic duality should be more than just a claim of a bulk geometry; it should also determine the matter fields and particles which can propagate in the bulk as well as the interactions. Ideally a holographic description of our own universe would include the standard model, neutrino masses, and other features of particle physics. We have taken up the challenge of determining the bulk theory for a particular example: the holographic duality between double-scaled SYK and JT-de Sitter space. The conjecture that $\rm DSSYK_{\infty}$} is dual to JT-de Sitter is not new but the identification of the bulk matter sector is. We’ve given evidence that the bulk theory is the ultra-strongly coupled fixed $\bar{g}$ limit of $(1+1)$ dimensional QCD—a string theory with mesons lying on a single Regge trajectory. Unlike the limit ’t Hooft limit with $\bar{\alpha}$ held fixed, the fixed $\bar{g}$ limit is a theory of interacting meson-strings with $g_{st}=\bar{g}$ being the open string coupling constant. On the $\rm DSSYK_{\infty}$} side the role of $g_{st}^{2}$ is played by $\lambda.$ The limit $\lambda\to 0$ corresponds to the standard ’t Hooft limit, in which mesons are non-interacting.